# Transforming Strategic Games into Biform Games: Applications in Allocation Mechanisms and Green Technology Investment

### v1

@@ 4.1 Problem Formulation
This section still uses the dual-game method in Section LABEL:Sec1 to study the issue of corporate green technology investment in market competition. It mainly focuses on two types of market competition models: The first type is the price competition model, specifically the Bertrand duopoly price competition model. The specific approach is to incorporate green technology investment items into the classic Bertrand model, establish a Bertrand model that includes green technology investment items, introduce cooperative pricing to construct a dual-game model, and interpret the famous Bertrand paradox through the comparison and analysis of non-cooperative game solutions and dual-game solutions. Meanwhile, it also proposes a specific path to resolve the price war dilemma and achieve win-win cooperation. For the second type, which targets the food supply chain, the Cournot model with green technology investment items is established, and cost sharing (or a cost-sharing mechanism) is introduced to construct a dual-type game model. Through discussion and analysis of the dual-type game solutions, the results show that enterprises in the supply chain can collaborate to increase corporate profits, and this outcome is fully consistent with raising the level of corporate green technology investment. It should be noted that in the above two types of models, improving the level of green technology investment is a requirement for maximizing corporate profits, and this is also the endogenous driving force for corporate green technology investment.
@@ 4.2.1 Problem Description
The Bertrand model is widely used to study the competitive decision-making behaviors of firms across various industries. For example, Ye et al. [Ye2008], Yang et al. [30], and Du Chuanzhong et al. [31] have explored the mutual influence and outcomes of decisions such as market pricing and promotional subsidies. Wang Aihu et al. [32] applied a non-cooperative-cooperative game framework to study the optimal decision-making of firms and governments in a mixed duopoly market. Zhang Wei et al. [33], Chen Qi and Wang Qiang [34], Kong Yue et al. [35], and Meng Zhan et al. [36] have used the Cournot and Bertrand models to analyze corporate research and development investment, especially within the context of corporate competition mechanisms and government-business interactions in green technological innovation, including the impact and driving forces of government subsidies.
@@
Considering the reality of both competition and cooperation in the market, this section will use the dual-type game model from Section LABEL:Sec1 to examine this issue. In fact, as long as there is a market, price acts as the ’invisible hand,’ driving market economic activities. The modern market economy differs from a laissez-faire economy: in addition to the market’s strong self-regulation function, there are also macroeconomic controls and even interventions. In terms of corporate price competition, aside from the dilemmas and deadlocks revealed by the Nash equilibrium in price wars, there are also cooperative, win-win strategies or tacit understandings aimed at survival and development. Reality shows that firms do not always strictly follow the Nash equilibrium of the Bertrand model; the ’killing a thousand, losing eight hundred’ decision in price wars is often avoided by both parties. Competition is undoubtedly the inevitable choice of individual rationality, but avoiding a lose-lose situation is also a unique wisdom of humanity.
@@ 4.4 The solution to the Biform game
where $\theta\in[0,1],\ \beta_{1},\beta_{2}\in\left[0,1\right],\ p\in I$. (4.4) is a Biform game derived from the strategic game ($1$), and is denoted as $(\left\{S,P,R\right\},v,X)$.
@@
Additionally, it should be noted that $\beta_{2}-l_{1}>0$¡ªi.e., $\beta_{2}>l_{1}$ and $l_{1}<1$¡ªand then we have
@@ 4.5 Summary of This Section
In this section, the Biform game analysis method established in Section LABEL:Sec1 is applied to study the issue of green technology investment by food enterprises in market competition. The research is mainly conducted based on two models: the Bertrand model and the supply chain Cournot model.
@@
The second part is based on the food supply chain model. A green technology investment term is introduced into this model, and cooperation is achieved by means of cost sharing of green technology this not only increases the level of green technology investment by enterprises within the chain, but also boosts their profits. This part mainly focuses on a food industry supply chain consisting of one supplier, one manufacturer, and one retailer. Considering the competition between the supply chains products and similar or substitute products in actual operations, both the selling price and green technology investment exert an impact on product demand, and also affect the respective profits of enterprises. Thus, to maximize profits, supply chain members compete with each other over the selling prices of their respective products and have respective demands for the manufacturer’s green technology investment leading to a non-cooperative scenario for this issue. In addition, members within the chain can also cooperate by sharing the manufacturer’s green technology investment costs, which means the issue also contains cooperative elements. In fact, the analysis results show that cooperation can increase the total profit of the supply chain. Therefore, for the supply chain Cournot model, we apply the Biform game model from Section LABEL:Sec1 to the study of this issue: we first establish a strategic game model, and then derive a Biform game from it. During the cooperation phase, through the supplier and retailer’s cost sharing of the manufacturer’s green technology investment, the level of green technology investment is increased, the total profit of the supply chain is raised, and Pareto optimality in the profits of members within the chain is achieved. On this basis, we analyzed the relationships between the total profit of the supply chain, the optimal price, and relevant parameters.

### v2

@@ 4.1 Problem Formulation
This section still uses the dual-game method in Section 2 to study the issue of corporate green technology investment in market competition. It mainly focuses on two types of market competition models: The first type is the price competition model, specifically the Bertrand duopoly price competition model. The specific approach is to incorporate green technology investment items into the classic Bertrand model, establish a Bertrand model that includes green technology investment items, introduce cooperative pricing to construct a dual-game model, and interpret the famous Bertrand paradox through the comparison and analysis of non-cooperative game solutions and dual-game solutions. Meanwhile, it also proposes a specific path to resolve the price war dilemma and achieve win-win cooperation. For the second type, which targets the food supply chain, the Cournot model with green technology investment items is established, and cost sharing (or a cost-sharing mechanism) is introduced to construct a dual-type game model. Through discussion and analysis of the dual-type game solutions, the results show that enterprises in the supply chain can collaborate to increase corporate profits, and this outcome is fully consistent with raising the level of corporate green technology investment. It should be noted that in the above two types of models, improving the level of green technology investment is a requirement for maximizing corporate profits, and this is also the endogenous driving force for corporate green technology investment.
@@ 4.2.1 Problem Description
The Bertrand model is widely used to study the competitive decision-making behaviors of firms across various industries. For example, Ye et al. [37], Yang et al. [38], and Du Chuanzhong et al. [39] have explored the mutual influence and outcomes of decisions such as market pricing and promotional subsidies. Wang Aihu et al. [40] applied a non-cooperative-cooperative game framework to study the optimal decision-making of firms and governments in a mixed duopoly market. Zhang Wei et al. [41], Chen Qi and Wang Qiang [42], Kong Yue et al. [43], and Meng Zhan et al. [44] have used the Cournot and Bertrand models to analyze corporate research and development investment, especially within the context of corporate competition mechanisms and government-business interactions in green technological innovation, including the impact and driving forces of government subsidies.
@@
Considering the reality of both competition and cooperation in the market, this section will use the dual-type game model from Section 2 to examine this issue. In fact, as long as there is a market, price acts as the ’invisible hand,’ driving market economic activities. The modern market economy differs from a laissez-faire economy: in addition to the market’s strong self-regulation function, there are also macroeconomic controls and even interventions. In terms of corporate price competition, aside from the dilemmas and deadlocks revealed by the Nash equilibrium in price wars, there are also cooperative, win-win strategies or tacit understandings aimed at survival and development. Reality shows that firms do not always strictly follow the Nash equilibrium of the Bertrand model; the ’killing a thousand, losing eight hundred’ decision in price wars is often avoided by both parties. Competition is undoubtedly the inevitable choice of individual rationality, but avoiding a lose-lose situation is also a unique wisdom of humanity.
@@ 4.4 The solution to the Biform game
where $\theta\in[0,1],\ \beta_{1},\beta_{2}\in\left[0,1\right],\ p\in I$. (4.4) is a Biform game derived from the strategic game (17), and is denoted as $(\left\{S,P,R\right\},v,X)$.
@@
Additionally, it should be noted that $\beta_{2}-l_{1}>0$ i.e., $\beta_{2}>l_{1}$ and $l_{1}<1$¡ªand then we have
@@ 4.5 Summary of This Section
In this section, the Biform game analysis method established in Section 2 is applied to study the issue of green technology investment by food enterprises in market competition. The research is mainly conducted based on two models: the Bertrand model and the supply chain Cournot model.
@@
The second part is based on the food supply chain model. A green technology investment term is introduced into this model, and cooperation is achieved by means of cost sharing of green technology this not only increases the level of green technology investment by enterprises within the chain, but also boosts their profits. This part mainly focuses on a food industry supply chain consisting of one supplier, one manufacturer, and one retailer. Considering the competition between the supply chains products and similar or substitute products in actual operations, both the selling price and green technology investment exert an impact on product demand, and also affect the respective profits of enterprises. Thus, to maximize profits, supply chain members compete with each other over the selling prices of their respective products and have respective demands for the manufacturer’s green technology investment leading to a non-cooperative scenario for this issue. In addition, members within the chain can also cooperate by sharing the manufacturer’s green technology investment costs, which means the issue also contains cooperative elements. In fact, the analysis results show that cooperation can increase the total profit of the supply chain. Therefore, for the supply chain Cournot model, we apply the Biform game model from Section 2 to the study of this issue: we first establish a strategic game model, and then derive a Biform game from it. During the cooperation phase, through the supplier and retailer’s cost sharing of the manufacturer’s green technology investment, the level of green technology investment is increased, the total profit of the supply chain is raised, and Pareto optimality in the profits of members within the chain is achieved. On this basis, we analyzed the relationships between the total profit of the supply chain, the optimal price, and relevant parameters.